cs161: Data Structures and Algorithms Handout #17 Michael Goldwasser Stanford University Tuesday, August 13, 1996

Final Exam Information

When: Saturday, 17 August, 1996, 12:15–3:15 P.M.

Where: Skiling 191

Crib sheet: The exam is closed book. You may, however, bring one 8\frac{1}{2}′′ × 11′′ sheet of paper as a crib sheet. Preparing a crib sheet can be a useful study aid, so take time in selecting material for it. You may use both sides of the paper and write as small as you like, but in any case, you are allowed only one sheet. Your must hand in your crib sheet when you hand in your exam.

Material covered: You are responsible for material covered in lectures, recitations, and problem sets from the entire course. You are also responsible for the material covered in the corresponding sections in the textbook.

Exam Conflict: If you are unable to attend the exam at the scheduled time, please talk to the TA in advance. There will be very little leniency without prior agreement.